# NavCoT: Boosting LLM-Based Vision-and-Language Navigation via Learning Disentangled Reasoning

### v1

@@ Abstract
Vision-and-Language Navigation (VLN), as a crucial research problem of Embodied AI, requires an embodied agent to navigate through complex 3D environments following natural language instructions. Recent research has highlighted the promising capacity of large language models (LLMs) in VLN by improving navigational reasoning accuracy and interpretability. However, their predominant use in an offline manner usually suffers from substantial domain gap between the VLN task and the LLM training corpus. This paper introduces a novel strategy called Navigational Chain-of-Thought (NavCoT), where we fulfill parameter-efficient in-domain training to enable self-guided navigational decision, leading to a significant mitigation of the domain gap in a cost-effective manner. Specifically, at each timestep, the LLM is prompted to forecast the navigational chain-of-thought by: 1) acting as a world model to imagine the next observation according to the instruction, 2) selecting the candidate observation that best aligns with the imagination, and 3) determining the action based on the reasoning from the prior steps. In this way, the action prediction can be effectively simplified benefiting from the disentangled reasoning. Through constructing formalized labels for training, the LLM can learn to generate desired and reasonable chain-of-thought outputs for improving the action decision. Experimental results across various training settings and popular VLN benchmarks (e.g., Room-to-Room (R2R), Room-across-Room (RxR), Room-for-Room (R4R)) show the significant superiority of NavCoT over the direct action prediction variants. Through simple parameter-efficient finetuning, our NavCoT outperforms a recent GPT4-based approach with $\sim$7relative improvement on the R2R dataset. We believe that NavCoT will help unlock more task-adaptive and scalable LLM-based embodied agents, which are helpful for developing real-world robotics applications. Code is available at https://github.com/expectorlin/NavCoT.
@@ A Introduction
In this paper, we propose Navigational Chain-of-Thought (NavCoT), where we conduct parameter-efficient in-domain training to enable LLMs to perform self-guided navigational reasoning for facilitating action decisions. Inspired by the world model theory [17, 18], when humans interact with the world, we tend to build a mental model that summarizes the surroundings we have seen before and helps us to predict the future. Then, we can make action decisions sequentially to complete different tasks based on this mental model. Therefore, we adapt the above process to the Chain-of-Thought (CoT) [19] reasoning mechanism in a trainable manner. The resulting strategy, termed navigational chain-of-thought, transforms the LLM into both a world model and a navigation reasoning agent, i.e., the LLM learns to imagine the future surroundings, filter the confusing observations based on the imagination, and then make final action decisions at each navigation timestep with the customized chain-of-thought labels. As shown in Fig. 1, through NavCoT, the LLM is able to filter redundant visual information based on the imagination and therefore significantly simplify the action decision.
@@ B.1 Vision-Language Navigation
VLN has received great attention and many works have been proposed in the past few years. Early approaches mainly focus on exploring data augmentation techniques [21, 22, 23, 24] and useful model architectures [25, 26, 27, 28] to alleviate data scarcity and improve cross-modal alignment. To further improve the generalization to unseen environments, pretraining-based approaches [29, 30, 31, 32, 33, 34, 35] have been widely developed in the VLN field. However, adapting to realistic application scenarios that require rich commonsense knowledge is still quite challenging for existing VLN agents. Furthermore, previous methods usually lack enough interpretability in action decisions. To address the above issues, a few recent works introduce LLMs with great knowledge storage as the navigation backbone in a zero-shot manner [14, 15]. Nevertheless, the severe domain gap and the dependency on high-cost LLMs significantly harm navigation performance and scalability.
This paper proposes a new LLM-based VLN approach called NavCoT, where we conduct parameter-efficient in-domain training for teaching LLMs to perform self-guided navigational reasoning for action decisions. Through NavCoT, the navigation performance and interpretability of the action prediction can be greatly enhanced cost-effectively.
@@ D Method
The overview of NavCoT is presented in Fig. 2. At each timestep $t$, a vision-to-text system is first employed to convert the surrounding observation into linguistic representations (Sec. D.1). Then, the LLM is prompted with the in-context example and textual represented navigation input to produce the navigational chain-of-thought (Sec. D.2). We adopt various training settings such as imitation-learning based finetuning and task-decomposed pretraining by collecting customized labels (Sec. D.3) to enable the LLM to generate reasonable and constrained outputs for improving the accuracy of action decisions (Sec. D.4).
@@ D.2 Navigational Chain-of-Thought Prompt
Since different tasks require distinguishing reasoning ability, proper design of the intermediate reasoning steps is crucial in designing the CoT prompt, which may greatly affect the performance of LLM’s predictions. In this work, we aim to empower the LLM to generate two significant intermediate reasoning steps for guiding the navigation action predictions, inspired by the world model theory [17, 18].
Concretely, the first reasoning step Future Imagination (FI) is to imagine the next observation according to the instruction and the navigation history. With FI, the LLM can monitor the progress of the navigation to guide itself for the subsequent action prediction. The second reasoning step, Visual Information Filter (VIF), aims to select the candidate observation that best aligns with the imagination generated in FI. Then the LLM generates the action prediction in the final Action Prediction (AP) reasoning step.
@@
Visual Information Filter (VIF). After generating the imagination in FI, we introduce a further reasoning step, VIF, to force LLM to explicitly select the observation that best matches the imagination from the redundant observation information. Through this explicit visual information filter procedure, the LLM can better learn to connect the imagination and action decision. Denote the option of observation that LLM predicts to align the imagination $U_{t}$ best as $V_{t}$. The desired output format for VIF is:
@@
D.3 Ground-Truth Collection
As shown in Fig. 3, due to the uncertainty of the LLM’s output and the complexity of the VLN task, it is hard for the LLM to generate the multi-step reasoning accurately for action decisions in a zero-shot manner. To address this issue, we collect the ground-truth of the navigational chain-of-thought based on the available VLN data for implementing in-domain training to improve the action decision.
We firstly collect the the ground-truth imagination ${U}^{*}_{t}$ for the reasoning task FI. Ideally, the ground-truth imagination should be consistent with the object/scene appearing in the following observation, which corresponds to the ground-truth action. Moreover, to enable better cross-modal alignment between the observation and the instruction for action decisions, the imagination is desired to be one of the mentioned objects/scenes in the given instruction. To this end, we resort to an LLM to extract the mentioned objects/scenes in the instruction and a cross-modal large model CLIP [50] to collect the ground-truth imagination ${U}^{*}_{t}$ at different timesteps. Concretely, for a given instruction $I$, we provide the prompt with the customized task example to ask the LLM to extract the mentioned objects/scenes from $I$. Denote the extracted landmark list from $I$ as $U^{la}=\{U^{la}_{k}\}_{k=1}^{M}$, where $M$ is the number of objects/scenes mentioned in the instruction $I$. For the ground-truth observation $B^{*}_{t}$ at each timestep $t$, we calculate the similarity between $B^{*}_{t}$ and each landmark $U^{la}_{k}$ in the list $U^{la}$, and take the one with the highest similarity as the ground-truth imagination label ${U}^{*}_{t}$:
@@ D.3 Ground-Truth Collection
D.4 In-domain Training
We conduct two kinds of in-domain training schemes in NavCoT, i.e., pretraining and finetuning, which are illustrated in the following.
@@ D.4 In-domain Training
Finetuning. Besides pretraining, we also conduct finetuning with the imitation learning strategy [1] to further adapt LLM to generate sequential action decisions for navigation. To this end, we reformulate the expert trajectory from original VLN dataset using $\mathrm{CoT}_{t}^{*}$ to construct instruction-following dataset. At each timestep $t$, we train the LLM to generate the complete navigational chain-of-thought $\mathrm{CoT}_{t}$ for action decisions. The finetuning objective $\mathcal{L}_{f}$ is:
$\mathcal{L}_{f}=-\sum_{t}\mathrm{CoT}^{*}_{t}\mathrm{log}(p_{\mathrm{LLM}}(\mathrm{CoT_{t}}|I,H_{t},D_{t})),$ (8)
After in-domain training, we prompt the trained LLMs to generate the navigational chain-of-thought for action decision based on the prompt and the in-context example described in Sec. D.2.
@@ E.1 Experimental Setup
Evaluation Metrics. The following standard metrics are used for evaluation on R2R [1] and REVERIE [2]:: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the target viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPL) [1]: success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible. Three evaluation metrics related to the instruction following are added for R4R [4] and RxR [5], i.e., the Coverage weighted by Length Score (CLS) [4], the normalized Dynamic Time Warping (nDTW) [52], and the Success weighted by nDTW (SDTW) [52].
Implementation Details. We train LLaMA-Adapter [20] and LLaMA 2 [8] of size 7B with 1.2M and 1.6M trainable parameters, respectively. To speed up the training, we use 4 V100 GPUs with a batch size of 8. The total training time lasts $\sim$10h on 4 V100 GPUs. The inference is conducted on a single V100 GPU. We use the AdamW optimizer with the learning rate of 0.001 and the weight decay of 0.02. For fast evaluation on various ablation experiments, we randomly choose 90 instruction-trajectory pairs from 8 scans in the val unseen split as the Val Unseen Subset. This subset serves as an efficient testbed to indicate the performance gap among different methods.
For collecting the the ground-truth imagination $U_{t}^{*}$, we utilize a powerful LLM tool, ChatGPT [53], to extract the landmarks from a given instruction. An in-context example for the landmark extraction is as follows:
Instruction: Walk along the rug past the statue on the wooden table.
Landmarks:
1.rug;
2.statue;
3.wooden table.
The maximum length of tokens in training and inference are set to 400 and 512, respectively. During inference, the temperature of LLM is set as 0. For parameter-efficient in-domain training of LLaMA 2 [8], we initialize the model weights using the pretrained LLaMA 2 weights with bias tuning [^1] https://llama2-accessory.readthedocs.io/en/latest/finetune/sg_peft.html#bias-norm-tuning-of-llama2-7b-on-alpaca. The training epoch for NavCoT based on both LLaMA-Adapter [20] and LLaMA 2 [8] is set as 2.
@@ E.2.1 Comparison with Existing Approaches
Table I presents the comparison between NavCoT and methods with different backbones on R2R [1]. From Table I, we can see that NavCoT with LLaMA 2 significantly outperforms a recent GPT4-based approach NavGPT [14] (e.g., 4.17% performance improvement on SPL), demonstrating that our method can effectively boost the performance of smaller LLMs (i.e., LLaMA 2) with a low training cost to outperform high-cost LLMs. We can also find that NavCoT with LLaMA-Adapter as backbone still shows a large performance gap with NavGPT, revealing the significant challenge of adapting LLMs with relatively small model capacity to the VLN task. Moreover, we can find that by adding a small amount of augmentation data (similar scale to that of R2R training data), the performance can be further enhanced largely, demonstrating the potential of improving our NavCoT via simple data augmentation.
Discussion. From the comparison results we can find that LLM-based agents like NavGPT [14] (using GPT-4) encounter a significant performance drop in VLN, demonstrating that original ability of LLM is far from solving such complex tasks. Our NavCoT proposes a promising and low-cost way to adapt smaller LLMs to VLN, which largely outperforms NavGPT. Note that although the language only backbone still shows a significant discrepancy with cross-modal backbone on different datasets, it provides much better interpretability through performing explicit reasoning than the latter, which is crucial in developing realistic robotic applications. In the future, it is promising to introduce our NavCoT into large vision-language models (VLMs) to further boost the navigation performance while enhancing interpretability.
@@ E.2.2 Ablation Study
We conduct extensive ablation experiments on R2R [1] to analyze the effect of each component in NavCoT, including training settings, reasoning tasks, navigation histories, etc. Training settings. Fig. 4 presents a comprehensive comparison among various training settings. From Fig. 4, we can draw three crucial conclusions: 1) NavCoT surpasses the DAP variant in all training settings, highlighting the power of explicit disentangled reasoning; 2) The great superiority of training-based settings over zero-shot ones shows the effectiveness of our parameter-efficient in-domain training strategy; 3) In-domain pretraining and finetuning can encourage the LLM to select actions that are relevant to the target position, reflected in significant improvement of OSR. However, the finetuning contributes much higher to the enhancement of SR and SPL than pretraining by benefiting from the learning of sequential action decisions.
@@
OSR SR SPL OSR SR SPL
@@
Setting Val Unseen
TL NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Example 1 10.34 6.91 44.10 34.53 31.00
Example 2 10.09 6.76 44.40 37.04 33.65
Example 3 10.34 6.74 44.83 36.14 32.50
Original Example 9.83 6.67 43.98 36.40 33.17
TABLE III: Ablation results for example sensitivity. The backbone is LLaMA 2 [8].
Setting Method Val Unseen Subset
NE $\downarrow$ SR $\uparrow$ SPL $\uparrow$
Finetuning No 9.67 13.33 5.25
All 7.45 21.11 14.55
Last 7.16 28.89 26.59
Pretraining then Finetuning No 7.92 22.22 8.21
All 7.13 16.67 11.26
Last 6.74 30.00 28.23
TABLE IV: Ablation results for navigation history. The backbone is LLaMA-Adapter [20].
@@
$\triangleright$ Vision-to-text system:
Coarse Direction & Object 7.23 33.33 26.67 24.20
Fine Direction & Caption 8.78 41.11 17.78 12.26
Coarse Direction & Caption (ours) 5.38 58.89 53.33 48.69
@@
$\triangleright$ Imagination formats:
Free Object 8.35 7.78 7.78 7.32
Direction & Object 5.99 53.33 44.44 41.12
Object (ours) 5.38 58.89 53.33 48.69
@@
$\triangleright$ In-context examples :
No-example training 7.97 54.44 20.00 13.62
NavCoT (ours) 5.38 58.89 53.33 48.69
TABLE V: Ablation results for vision-to-text systems, chain-of-thought prompts, imagination formats, backbones, and in-context examples.* denotes the zero-shot manner.
@@
Example sensitivity. We test NavCoT on different in-context examples for in-domain training to examine the example sensitivity of NavCoT. Concretely, we define three more examples, using different manually written instructions and observation descriptions. From the results in Table III, we can see that NavCoT encounters a small range of performance changes when given different in-context examples, showing low example sensitivity of NavCoT.
Navigation history. We also conduct experiments to verify how navigation history influences navigational reasoning, as shown in Table IV. In our implementation, “No” means removing history information in the input. “All” and “Last” represent preserving all previous steps and only the last step as history, respectively. From Table IV, we can find that “No” performs the worst among all settings not surprisingly, since the navigation history is essential for monitoring the progress of navigation. Moreover, the “Last” setting performs best among all settings. A possible reason is that the “All” setting may bring much more noise and also exceed the maximum sequence length, therefore leading to performance degradation.
Other ablations. Table V shows the ablation results for vision-to-text systems, chain-of-thought prompts, imagination formats, backbones, and in-context examples. In Table V: 1) Coarse Direction means the textual direction description does not contain the concrete angle information (see Sec. D.1), and Fine Direction means adding the angle information into the textual direction description; 2) For Coarse Direction & Object setting, we use the GroundingDino [54] model combined with the RAM [55] model to extract the object information in the observation; 3) For original CoT, we directly prompt the LLM to generate thought and action like [14]; 4) For Free Object setting, we extract the object of ground-truth action (observation) rather than using the object mentioned in the instruction as the imagination label.
We can draw the following conclusions from Table V: 1) vision-to-text system: By comparing Coarse Direction & Object and Coarse Direction & Caption (ours), we can find that captions are more suitable for our NavCoT, probably because captions contain spatial information for facilitating the alignment to the instruction, while pure object detections may have more noise. Through the comparison between Fine Direction and Coarse Direction, we can also find that fine-grained direction may bring unnecessary noise for confusing the navigation decision. This is reasonable since instructions generally contain coarse direction descriptions such as “go up” and “turn left”. 2) Chain-of-Thought prompts: We can see that NavCoT outperforms the original CoT, suggesting that our design of combining the world model into CoT can better activate reasonable navigational reasoning. Benefiting from training with formalized labels, NavCoT outperforms both the zero-shot variant and the original CoT for GPT-4 [16] largely (e.g., $\sim$43 and $\sim$36 points in SPL, respectively). Moreover, note that our NavCoT also shows great superiority than the original CoT for GPT-4 [16] in the computational efficiency, i.e., the inference time of one-step decision for GPT-4 and NavCoT are $\sim$9.8s and $\sim$0.5s, respectively. 3) Imagination formats: the great superiority of Object (ours) over Free Object reveals the reasonability and the effectiveness of our NavCoT by constraining the imagination to be mentioned objects in the instruction. This strategy essentially fulfills a more task-oriented world model for better instruction following and modality alignment. Moreover, asking the LLM to imagine both the direction and object brings performance degradation, probably because the imagined direction information is usually inaccurate and harms the alignment with the instruction, and therefore confuses the LLM for action decisions. 4) Backbones: we can find that NavCoT+LLaMA-Adapter is largely inferior to its alternative NavCoT+LLaMA2. This is not surprising since LLaMA-Adapter is based on LLaMA 1 [7] while LLaMA 2 is trained on 40% more data than the previous version. Moreover, LLaMA-Adapter has 0.4M fewer trainable parameters than the bias tuning setting [49]. However, NavCoT shows consistent improvement on both backbones under our in-domain training strategies. Interestingly, we can find that LLaMA 2 outperforms GPT-4 in the zero-shot NavCoT setting, probably because when facing the prompt with less information and constrained output format, the advantage of GPT-4 is not obvious over the smaller language model LLaMA 2. However, such kind of prompt is more beneficial for implementing in-domain training. 5) In-context examples: the comparison between No-example training and NavCoT shows that adding the example is still necessary under the training setting and the in-context example can effectively facilitate the training.
@@
SR SPL CLS nDTW SDTW
@@
TABLE VI: Comparison results on RxR English subset. * denotes our re-implementation results with the imitation learning strategy [30]. The best results for cross-modal and language-only backbones are denoted by bold and blue fonts, respectively.
@@
TABLE VII: Comparison results on REVERIE. * denotes our re-implementation results with the imitation learning strategy [30]. The best results for cross-modal and language-only backbones are denoted by bold and blue fonts, respectively.

### v2

@@ Abstract
Vision-and-Language Navigation (VLN), as a crucial research problem of Embodied AI, requires an embodied agent to navigate through complex 3D environments following natural language instructions. Recent research has highlighted the promising capacity of large language models (LLMs) in VLN by improving navigational reasoning accuracy and interpretability. However, their predominant use in an offline manner usually suffers from substantial domain gap between the VLN task and the LLM training corpus. This paper proposes a novel strategy called Navigational Chain-of-Thought (NavCoT), where we fulfill parameter-efficient in-domain training to enable self-guided navigational decision, leading to a significant mitigation of the domain gap in a cost-effective manner. Specifically, at each timestep, the LLM is prompted to forecast the navigational chain-of-thought by: 1) acting as a world model to imagine the next observation according to the instruction, 2) selecting the candidate observation that best aligns with the imagination, and 3) determining the action based on the reasoning from the prior steps. In this way, the action prediction can be effectively simplified benefiting from the disentangled reasoning. Through constructing formalized labels for training, the LLM can learn to generate desired and reasonable chain-of-thought outputs for improving the action decision. Experimental results across various training settings and popular VLN benchmarks (e.g., Room-to-Room (R2R), Room-across-Room (RxR), Room-for-Room (R4R)) show the significant superiority of NavCoT over the direct action prediction variants. Through simple parameter-efficient finetuning, our NavCoT outperforms a recent GPT4-based approach with $\sim$7relative improvement on the R2R dataset. We believe that NavCoT will help unlock more task-adaptive and scalable LLM-based embodied agents, which are helpful for developing real-world robotics applications. Code is available at https://github.com/expectorlin/NavCoT.
@@ A Introduction
In this paper, we propose Navigational Chain-of-Thought (NavCoT), which enables LLMs to learn to perform disentangled navigational reasoning powered by parameter-efficient in-domain training. Inspired by the world model theory [17, 18], when humans interact with the world, we tend to build a mental model that summarizes the surroundings we have seen before and helps us to predict the future. Then, we can make action decisions sequentially to complete different tasks based on this mental model. Therefore, we adapt the above process to the Chain-of-Thought (CoT) [19] reasoning mechanism in a trainable manner. The resulting strategy, termed navigational chain-of-thought, transforms the LLM into both a world model and a navigation reasoning agent, i.e., the LLM learns to imagine the future surroundings, filter the confusing observations based on the imagination, and then make final action decisions at each navigation timestep with the customized chain-of-thought labels. As shown in Fig. 1, through NavCoT, the LLM is able to filter redundant visual information based on the imagination and therefore significantly simplify the action decision.
@@ B.1 Vision-Language Navigation
VLN has received great attention and many works have been proposed in the past few years. Early approaches mainly focus on exploring data augmentation techniques [21, 22, 23, 24], introducing learning mechanisms [25, 26, 27, 28], and designing model architectures [29, 30, 31, 32] to alleviate data scarcity and enhance the navigation performance. To further improve the generalization to unseen environments, pretraining-based approaches [33, 34, 35, 36, 37, 38, 39] have been widely developed in the VLN field. However, adapting to realistic application scenarios that require rich commonsense knowledge is still quite challenging for existing VLN agents. Furthermore, previous methods usually lack enough interpretability in action decisions. To address the above issues, a few recent works introduce LLMs with great knowledge storage as the navigation backbone in a zero-shot manner [14, 15, 40, 41]. NavGPT [14] constructs a purely LLM-based navigation agent that receives textual represented observations and navigation histories to make action decisions. DiscussNav [15] incorporates large models with unique capabilities to act as domain experts and asks the agent to actively discuss with these experts to make decisions. MapGPT [40] uses an online constructed language-formalized map to encourage the agent for global exploration and adaptive planning. Nevertheless, the severe domain gap and the dependency on high-cost LLMs significantly harm navigation performance and scalability. NaviLLM [42] constructs a trainable LLM-based agent to make LLMs to better adapt to VLN tasks. However, it directly maps navigation inputs to action predictions without the reasoning output.
This paper proposes a new LLM-based VLN approach called NavCoT, where we conduct parameter-efficient in-domain training for teaching LLMs to perform disentangled navigational reasoning accurately for facilitating action decisions. Our NavCoT significantly improves the scalability of LLM-based VLN agents and narrows the domain gap between LLM’s training corpus and the VLN task. Through navigational chain-of-thought reasoning generation in a trainable manner, our NavCoT improves the accuracy of both navigation reasoning and action decisions.
@@ D Method
The overview of NavCoT is presented in Fig. 2. At each timestep $t$, we adopt a vision-to-text system to convert the surrounding observation into linguistic representations (Sec. D.1). Then, we prompt the LLM with the in-context example and textual represented navigation input to produce the disentangled navigational reasoning, i.e., navigational chain-of-thought (CoT) (Sec. D.2). We collect customized reasoning labels (Sec. D.3) and adopt various training settings such as imitation-learning based finetuning and task-decomposed pretraining (Sec. D.4) for implementing parameter-efficient in-domain training. As a result, the LLM learns to generate correct disentangled reasoning with constrained formats for improving the action decision accuracy.
@@ D.2 Navigational Chain-of-Thought Prompt
Proper design of the intermediate reasoning steps is crucial in designing the chain-of-thought (CoT) prompt, which may greatly affect the performance of LLM’s predictions. In this work, we aim to empower the LLM to generate two significant intermediate reasoning steps for guiding the navigation action predictions, inspired by the world model theory [17, 18]. Specifically, we formalize the navigational chain-of-thoughts to contain three important steps, i.e., Future Imagination (FI), Visual Information Filter (VIF), and Action Prediction (AP). FI enables the agent to predict possible objects/scenes it may encounter with the guidance of the instruction, which helps the agent capture the environmental dynamics and monitor the progress of the navigation. VIF explicitly connects the generated imagination with the subsequent action prediction.
@@
Visual Information Filter (VIF). After generating the imagination in FI, we introduce a further reasoning step, VIF, to force LLM to explicitly select the observation that best matches the imagination from the redundant observation information. Denote the option of observation that LLM predicts to align the imagination $U_{t}$ best as $V_{t}$. The desired output format for VIF is:
@@
D.3 Reasoning Ground-Truth Collection
As shown in Fig. 3, due to the uncertainty of the LLM’s output and the complexity of the VLN task, it is hard for the LLM to generate the multi-step reasoning accurately for action decisions in a zero-shot manner. For example, in Fig. 3(a), the LLM produces confusing reasoning output for the visual information filter. In Fig. 3(b), the LLM generates the wrong imagination “bed” that would mislead itself to choose the wrong action. To effectively adapt the LLM for complex navigation decisions in the VLN task, we collect the ground-truth of the navigational chain-of-thought based on the available VLN data for implementing in-domain training.
We firstly collect the ground-truth imagination ${U}^{*}_{t}$ for the reasoning task FI. Ideally, the ground-truth imagination should be consistent with the object/scene appearing in the following observation, which corresponds to the ground-truth action. Moreover, to enable better cross-modal alignment between the observation and the instruction for action decisions, the imagination is desired to be one of the mentioned objects/scenes in the given instruction. To this end, we resort to an LLM to extract the mentioned objects/scenes in the instruction and a cross-modal large model CLIP [57] to collect the ground-truth imagination ${U}^{*}_{t}$ at different timesteps. Concretely, for a given instruction $I$, we provide the prompt with the customized task example to ask the LLM to extract the mentioned objects/scenes from $I$. Denote the extracted landmark list from $I$ as $U^{la}=\{U^{la}_{k}\}_{k=1}^{M}$, where $M$ is the number of objects/scenes mentioned in the instruction $I$. For the ground-truth observation $B^{*}_{t}$ at each timestep $t$, we calculate the similarity between $B^{*}_{t}$ and each landmark $U^{la}_{k}$ in the list $U^{la}$, and take the one with the highest similarity as the ground-truth imagination label ${U}^{*}_{t}$:
@@ D.3 Reasoning Ground-Truth Collection
D.4 In-domain Chain-of-Thought Training
Pretraining and finetuning are two widely used techniques for training VLN agents [58, 34, 35]. Therefore, we also conduct both pretraining and finetuning with VLN data for fulfilling our in-domain training scheme. Specifically, we decompose each reasoning procedure in the navigational chain-of-thought (CoT) as a separate task to implement multi-task pretraining. As a result, the accuracy of each reasoning procedure can be effectively promoted for facilitating the subsequent finetuning. After pretraining, we conduct finetuning with the imitation learning strategy [1] to ask the LLM to produce complete navigation CoT reasoning, which further adapts LLM to generate self-guided navigational decisions sequentially for successful navigation.
@@ D.4 In-domain Chain-of-Thought Training
Finetuning. We reformulate the expert trajectory from the original VLN dataset using $\mathrm{CoT}_{t}^{*}$ to construct instruction-following finetuning dataset. At each timestep $t$, we train the LLM to generate the complete navigational chain-of-thought $\mathrm{CoT}_{t}$ for action decisions. The finetuning objective $\mathcal{L}_{f}$ is:
$\mathcal{L}_{f}=-\sum_{t}\mathrm{CoT}^{*}_{t}\mathrm{log}(p_{\mathrm{LLM}}(\mathrm{CoT_{t}}|I,H_{t},D_{t})).$ (8)
After in-domain training, we prompt the trained LLMs to generate the navigational chain-of-thought for action decision based on the prompt and the in-context example described in Sec. D.2. Benefiting from our introduced in-domain training strategy, the LLM can learn to generate disentangled navigational reasoning with the desired format to guide itself for accurate action decisions.
@@ E.1 Experimental Setup
Evaluation Metrics. The following standard metrics are used for evaluation on R2R [1] and REVERIE [2]: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the target viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPL) [1]: success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible. Three evaluation metrics related to the instruction following are added for R4R [4] and RxR [5], i.e., the Coverage weighted by Length Score (CLS) [4], the normalized Dynamic Time Warping (nDTW) [60], and the Success weighted by nDTW (SDTW) [60].
Implementation Details. We train LLaMA-Adapter [20] and LLaMA 2 [8] of size 7B with 1.2M and 1.6M trainable parameters, respectively. To speed up the training, we use 4 V100 GPUs with a batch size of 8. The total training time lasts $\sim$10h on 4 V100 GPUs. The inference is conducted on a single V100 GPU. We use the AdamW optimizer with the learning rate of 0.001 and the weight decay of 0.02. For fast evaluation on various ablation experiments, we randomly choose 90 instruction-trajectory pairs from 8 scans in the val unseen split as the Val Unseen Subset. This subset serves as an efficient testbed to indicate the performance gap among different methods. More implementation details are given in the supplementary material.
@@ E.2.1 Comparison with Existing Approaches
Table I presents the comparison between NavCoT and methods with different backbones on R2R [1]. From Table I, we can see that NavCoT with LLaMA-Adapter as backbone shows a large performance gap with NavGPT [14], revealing the significant challenge of adapting LLMs with relatively small model capacity to the VLN task. This shows LLM’s original capability is a key factor influencing the navigation performance. However, by plugging into another small language model, i.e., LLaMA2 7B [8], NavCoT outperforms NavGPT which employs GPT-4 as backbone in almost all metrics (4.17%, 2.4%, and 1.98% improvement in SPL, SR, and OSR, respectively). By adding a small amount of augmentation data (similar scale to that of R2R training data), NavCoT further surpasses NavGPT significantly in all metrics (7.64%, 6.23%, and 6.11% improvement in SPL, SR, and OSR, respectively). These results show the effectiveness of our NavCoT in adapting open-source affordable LLMs to the VLN task.
Discussion. Note that recent works using LLMs for improving the VLN task, e.g., NavGPT [14], DiscussNav [15], and MapGPT [40], generally show a performance drop compared with previous state-of-the-art VLN methods like scaleVLN [39]. This demonstrates that the original ability of LLM is far from solving complex navigation tasks. Nevertheless, LLM-based agents still have unique advantages for assisting the VLN task. Firstly, these GPT4-based approaches reveal the potential of zero-shot VLN which is more practical in real-world applications. Moreover, LLM-based agents can produce navigational reasoning output, which significantly improves decision interpretability and facilitates interaction with humans. These can not be achieved by previous navigation models.
Besides enhancing the interpretability and interactivity, our NavCoT introduces trainable reasoning, which adapts a relatively small LLM to the VLN task and shows superiority over GPT4-based methods in improving the accuracy of both navigational reasoning and action decisions. Our proposed method provides a scalable solution to mitigate the domain gap between the LLM’s training corpus and the VLN task. In the future, it is promising to introduce our NavCoT into large vision-language models (VLMs) to further boost the navigation performance while enhancing interpretability.
@@ E.2.2 Ablation Study
In this subsection, we present the ablation results, including training settings, reasoning tasks, chain-of-thought prompts, and backbones, to analyze the effect of each component in NavCoT. More results are given in the supplementary material.
Training settings. Fig. 4 presents a comprehensive comparison among various training settings. From Fig. 4, we can draw three crucial conclusions: 1) NavCoT surpasses the DAP variant in all training settings, highlighting the power of explicit disentangled reasoning; 2) The great superiority of training-based settings over zero-shot ones shows the effectiveness of our parameter-efficient in-domain training strategy; 3) In-domain pretraining and finetuning can encourage the LLM to select actions that are relevant to the target position, reflected in significant improvement of OSR. However, the finetuning contributes much higher to the enhancement of SR and SPL than pretraining by benefiting from the learning of sequential action decisions.
@@
OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$
@@
TABLE III: Ablation results for chain-of-thought prompts and backbones. * denotes the zero-shot manner.
@@
Chain-of-thought prompts. Table III shows the ablation results for chain-of-thought prompts, where we can see that NavCoT outperforms the original CoT, suggesting that our design of combining the world model into CoT can better activate reasonable navigational reasoning. Benefiting from training with formalized labels, NavCoT outperforms both the zero-shot variant and the original CoT for GPT-4 [16] largely (e.g., $\sim$43 and $\sim$36 points in SPL, respectively). Moreover, note that our NavCoT also shows great superiority than the original CoT for GPT-4 [16] in the computational efficiency, i.e., the inference time of one-step decision for GPT-4 and NavCoT are $\sim$9.8s and $\sim$0.5s, respectively.
Backbones. We can find in Table III that NavCoT+LLaMA-Adapter is largely inferior to its alternative NavCoT+LLaMA2. This is not surprising since LLaMA-Adapter is based on LLaMA 1 [7] while LLaMA 2 is trained on 40% more data than the previous version. Moreover, LLaMA-Adapter has 0.4M fewer trainable parameters than the bias tuning setting [56]. However, NavCoT shows consistent improvement on both backbones under our in-domain training strategies. Interestingly, we can find that LLaMA 2 outperforms GPT-4 in the zero-shot NavCoT setting, probably because when facing the prompt with less information and constrained output format, the advantage of GPT-4 is not obvious over the smaller language model LLaMA 2. However, such kind of prompt is more beneficial for implementing in-domain training.
We also present the result of using the imagination and CLIP-similarity to make action decisions, which we name as “LLM & CLIP*”. Concretely, we utilize the imagination, i.e., imagined landmark generated by NavCoT (LLaMA2) to select the observation (action) by calculating the CLIP-smilarity between the textual landmark feature and the observation feature. From Table III, we can find that LLM & CLIP* outperforms LLaMA 2*, LLaMA-Adapter, LLaMA-Adapter*, and GPT-4*, showing that using CLIP for matching the imagination (i.e., the textually represented landmark) with the right observations has relatively high accuracy for the action decision. This demonstrates the reasonability of our introduced strategy for collecting imagination ground-truths (GTs) using LLM and CLIP. We can also find that LLaMA 2 (ours) surpasses LLM & CLIP* by a large margin. This is reasonable since the navigation decision relies on comprehensive factors including landmarks, navigation history, direction, etc. This result further reveals the importance of our parameter-efficient in-domain training strategy that effectively adapts the LLM backbone for making complex navigation decisions.
@@
SR $\uparrow$ SPL $\uparrow$ CLS $\uparrow$ nDTW $\uparrow$ SDTW $\uparrow$
@@
TABLE IV: Comparison results on RxR English subset. * denotes our re-implementation results with the imitation learning strategy [34]. The best results for cross-modal and language-only backbones are denoted by bold and blue fonts, respectively.
@@
TABLE V: Comparison results on REVERIE. * denotes our re-implementation results with the imitation learning strategy [34]. The best results for cross-modal and language-only backbones are denoted by bold and blue fonts, respectively.